# A Distributionally-Robust Framework for Nuisance in Causal Effect Estimation

### v1

@@ 1 Introduction
Causal inference enables us to assess the impact of a treatment action. Its application originated in the policy-making field [23] including healthcare [32]. Recently, the focus was expanded to individualized decision-making such as precision medicine [32], recommendation [33, 6], and advertisement [36, 40] with the help of advanced machine learning-based methods.
@@
We estimate the causal effect of a treatment action (e.g., prescription of a specific medicine). That is, we need accurate predictions of both potential outcomes with and without the treatment to take its difference since the supervision of the actual effect itself is never given. In observational data, actions are not assigned randomly but selected by past decision-makers. Therefore, models must generalize beyond the factual data distribution to handle the systematically missing counterfactual outcomes. This is called the fundamental problem of causal inference [34].
@@
Conventional statistical methods for causal inference deal with this action selection bias by matching extraction or importance sampling [30]. A representative and versatile approach is inverse probability weighting using propensity scores (IPW) [3]. IPW estimates propensity scores—the probability of past treatment decisions—then weights instances by their inverse to account for selection bias. This two-step strategy has fundamental limitations: First, the final accuracy depends entirely on propensity score precision. Second, extreme weights can concentrate on few samples, reducing effective sample size and causing high estimation variance [20]. The overall accuracy can only be guaranteed asymptotically, limiting its applicability to modern non-asymptotic scenarios such as high dimensional models as DNNs for capturing complex heterogeneity or complex action spaces.
@@
Various countermeasures have been tried to alleviate this problem, such as doubly robust (DR) methods [20, 21, 13] and double machine learning [9, 26], which are clever combinations of outcome prediction models and only weighting its residuals using estimated propensity. Nevertheless, the IPW-based approach’s limitation is the instability of the two-step procedure in non-asymptotic scenarios. Its large estimation variance is problematic in cases of high dimensional covariates or cases where the propensity score is close to 00 or $1$ and thus the effective sample size is limited [2].
@@
As in various other fields, advancement with deep neural networks (DNN) has gained substantial attention in causal inference literature [24]. One of the notable advancements made when [34] and [18] introduced DNN to causal inference was representation-level balancing through distribution discrepancy measures. The representation extracted from the input covariates is encouraged to be balanced, i.e., independent of the action, by measuring and minimizing the discrepancy between the representation distributions conditioned on the action. Representation balancing provides non-asymptotic performance guarantees through generalization error bounds [34]. This results in an end-to-end training procedure, free from the concerns in the intermediate estimation problem. However, it has been pointed out that the guarantee of generalization error of representation balancing requires the unrealistic assumption of invertible representation extraction [19, 43, 42]. In recent years, researchers have explored approaches that combine weighting with representation learning [16, 41, 8]. However, to the best of our knowledge, this line of research has not yet yielded a theoretically grounded approach based on generalization error analysis.
@@
We analyze the generalization error of the weighted loss function using a propensity model, which is implemented as adversarial end-to-end learning of the target and weighting models. We define the worst-case loss with respect to the nuisance model ambiguity as a general framework. Our adversarial loss simultaneously accounts for the true propensity’s ambiguity and the statistical instability due to skewed weights. We demonstrate our framework’s effectiveness by applying it to two established methods: the doubly robust learner [21] and representation-based architectures [10]. Experiments on both synthetic and real-world datasets show consistent improvements over baseline methods.
@@ 2 Problem setting
We consider a standard causal inference framework. We have observational data $D=\{(x^{(n)},a^{(n)},y^{(n)})\}_{1}^{N}$ with $N$ i.i.d. instances. Each instance contains a $d$-dimensional background feature $x^{(n)}\in\mathcal{X}\subset\mathbb{R}^{d}$, a treatment action $a^{(n)}\in\mathcal{A}=\{0,1\}$, and an outcome $y^{(n)}\in\mathcal{Y}$. In the Neyman-Rubin potential outcome framework [31], the potential outcomes of both factual and counterfactual actions are expressed as random variables $\{Y_{a}\}_{a\in\{0,1\}}$, of which only the factual outcome is observed $\left(y^{(n)}=y_{a^{(n)}}\right)$ and the counterfactual one $y_{1-a^{(n)}}$ is missing.
@@
As a sufficient condition for consistent learnability of the CATE, we follow the standard set of assumptions in the potential outcome framework [17].
@@ 3 Related work
Orthogonal statistical learning, such as the DR-Learner [21] and the R-Learner [26], utilize two kinds of nuisance models of predicting outcomes and predicting treatments. These methods have been shown to be robust to estimation errors for the first-stage nuisance parameters of propensity and outcome models. Specifically, the errors do not affect the final estimation in the first-order sense of the Taylor expansion. However, their main limitation lies in high estimation variance when applied to non-asymptotic situations [2].
@@
Starting with [18, 34], a number of causal inference methods based on DNNs and representation balancing have been proposed [24]. The representation-based methods have been demonstrated to be superior in complex problems such as nonlinear responses [18], large actions spaces [38] including continuous [25] or structured spaces [14], and so forth. These are end-to-end methods based on adversarial formulations. They virtually evaluate the worst-case with respect to the uncertainty of the model by distribution discrepancy between the representations of covariates in treated and control groups. On the other hand, representation balancing has certain limitations in an unrealistic theoretical assumption that the representation extractor should be invertible. It is shown that unobservable error terms arise when the invertibility is violated [19, 43, 42].
A solution to this problem is the representation decomposition [16, 41, 8]. They aim to identify confounding factors that affect both action selection and the outcomes and weights with only those factors. Joint optimization approaches have also been proposed for ATE estimation [35]. However, their formulations have no guarantee as a joint optimization, e.g., in the form of generalization error bound for all possible combinations of the target and nuisance models. Even though estimation of weights with other model parameters (i.e., outcome model and representations) fixed is guaranteed, and vice versa, though joint optimization may lead to unexpected results. Weights should be optimized to balance the distributions, but especially when the noise is heterogeneous, lowering the weights to the noisier regions may help the (wrongly) weighted loss of the outcome prediction model. Thus, we aim at a principled and versatile weighting method while incorporating the advantages of end-to-end modeling by adversarial formulation.
@@
Our proposed method can be interpreted in the context of distributionally robust optimization (DRO). DRO aims to achieve robust learning against discrepancies between the empirical distribution of training data and the actual test distribution [28]. DRO is formalized as follows.
@@
Recent efforts in offline reinforcement learning revealed the benefit of pessimism on the candidate assessment [29, 7]. In reinforcement learning, we consider modeling the cumulative expected reward in the long run as the Q function for assessing each action at each time step. The Q function is supposed to be maximized with respect to action $a$ during the inference phase. If the estimation error on an action is optimistic, i.e., if the Q value is overestimated, the action is likely to be selected over other better candidates. Therefore, conservative modeling of Q-function is preferable [22], i.e., training a model to estimate below the true value when uncertain. The provable benefit of pessimism has been revealed in recent years [7]. We apply this pessimism principle to weighted estimation in causal inference; that is, our method pessimistically estimates the weighted loss with respect to the uncertainty of the weights. Our method applies the principle of pessimism in training that minimize the balanced loss with loss uncertainty due to propensity ambiguity and extreme weights, which can be evaluated by the generalization error analysis.
@@ 4.1 Adversarial Reformulation of Plug-in Nuisance
$\hat{\theta}$ $=\operatorname*{arg\,min}_{\theta\in\Theta}\hat{L}(\theta;\hat{\mu}).$ (4)
The assumed model classes are denoted by $M$ and $\Theta$. A typical form of 4 is instance weighted loss:
@@
Plug-in estimators typically analyze generalization error through stage-wise convergence rates [27]. For complex model classes, both convergence rates and their coefficients matter. We therefore adopt an end-to-end approach analyzing generalization error directly [34], avoiding two-step estimation issues.
@@
The second term (b) in Section 4.1 is upper-bounded by the excess risk of the weighted loss since $\mu^{*}\in M$.
[…]
@@
The proofs for these theorems can be found in Appendix 0.A.
Consistent with findings and proposals in several prior studies [4, 39, 37], our analysis also demonstrates that the mean squared weight is related to the stability of the loss (b) in Section 4.1. In other words, assigning too large weights to a small fraction of the sample compromises the effective sample size. Therefore, to mitigate the impact of (b), we introduce a regularization on $\mu$ in the form of the squared weights, albeit at the potential cost of increasing the misspecification error (a) in Section 4.1. More precisely, we define the class for restricted squared weights $\mathcal{R}$ as follows.
=$\mathcal{R}=\left\{\mu\in M_{0}\middle|\frac{1}{N}\sum_{n}^{N}\left(w^{(n)}_{\mu}\right)^{2}\leq C\right\},$ (13)
@@ 4.2 Nuisance-robust Doubly-robust Network (NuDRNet)
Under the aforementioned approach, next, we will discuss the application to specific estimation methods. Among two-step methods with weighting, we take the doubly-robust learner or DRNet [21] as a simple but clever baseline method. DRNet regresses a transformed target variable on $x$, which is calculated by the plug-in CATE estimate $\hat{f}_{1}(x^{(n)})-\hat{f}_{0}(x^{(n)})$ with residual adjusted with weights:
@@
We propose nuisance-robust DRNet (NuDRNet) as an extension of DRNet. Our approach starts with a pre-trained propensity $\hat{\mu}$ as the initial solution $\mu_{0}$ and a randomly initialized parameter $\theta$. We then perform adversarial optimization as described in 7, using the mean squared error: $\hat{L}(\theta;\mu)=\frac{1}{N}\sum_{n}\left(z^{(n)}_{\mu}-\tau_{\theta}(x^{(n)})\right)^{2}.$ Unlike DRNet, which uses a fixed pre-trained $\hat{\mu}$, NuDRNet perturbs $\hat{\mu}$ adversarially during training of $\hat{\tau}_{\theta}$.
Next, we discuss how to incorporate the constraint $\mu\in M$ into a gradient-based update. A typical implementation is as regularization term such as $\alpha\max\{0,\hat{E}(\mu)-c\},$ where $c$ is the tolerance parameter and is set to the evidence of the pre-trained solution with early stopping $c=\hat{E}(\mu_{0})$. However, since $-\hat{L}(\theta;\mu)$ is not convex with respect to $\mu$, the regularization cannot reproduce the constrained optimization within $\mathcal{U}$. To address this issue, we implement the augmented Lagrangian method [5] for handling the constraint $\mu\in\mathcal{U}$.
@@
In each epoch, we minimize $\hat{J}(\theta,\mu)$ with respect to $\theta$ and maximize with respect to $\mu$. Overall, our loss controls the error due to the uncertainty of $\mu$ (4.1-a) by maximizing the first term with respect to $\mu$ under the likelihood constraint in the third and the fourth terms, while simultaneously controlling the estimation variance of the weighted empirical loss (4.1-b) by flattening the weight with the second term. The parameters of evidence terms $\alpha_{k}$ and $\lambda_{k}$ are updated according to the augmented Lagrangian method. The whole algorithm is summarized in Algorithm 1 and Figure 1 illustrates the architecture.
@@
1:  Train $f_{1},f_{0},\mu$ by an arbitrary supervised learning method, e.g.: $\hat{f}_{a}\leftarrow\operatorname*{arg\,min}_{f_{a}}\frac{1}{N}\sum_{n:a^{(n)}=a}(y^{(n)}-f_{a}(x^{(n)}))^{2}$ for each $a\in\{0,1\}$,$\mu_{0}\leftarrow\operatorname*{arg\,min}_{\mu}-\frac{1}{N}\sum_{n}a^{(n)}\log\mu(x^{(n)})+(1-a^{(n)})\log(1-\mu(x^{(n)}))$
@@
6:        Update parameters with objective 19 and step sizes $\eta_{\theta},\eta_{\mu}$ from optimizes:
@@ 4.3 Nuisance-robust Shared Representation Network (NuSNet)
We further investigate exploiting representation-based architectures. Deep networks have also performed well in the context of causal inference owing to inductive bias of their representation layers. Among such methods, SNet [10] has shown outstanding performance thanks to its flexible architecture. SNet has three heads and five representation extractors. Specifically, there is a head for $y_{1}$, $y_{0}$, and $\mu$, respectively, and a confounder representation $\phi_{c}$ shared by the three heads, a representation shared by the two outcome heads $\phi_{o}$, and the representations unique to each hypothesis layer $\phi_{1},\phi_{0},\phi_{\mu}$. Denoting heads as $h_{\cdot}$, the SNet predictor is expressed as follows.
@@
The detailed architecture is illustrated in Figure 1(a).
We apply adversarial joint learning to this architecture. The original SNet does not employ weighted loss. As noted before, simply applying non-adversarial joint optimization of weight and the outcome model leads to catastrophic results, as seen in our experiment in Appendix 0.E. We therefore first pre-train as in the original SNet and then tune with an adversarially-weighted risk and other regularization terms. In the tuning phase, the shared representation layer is fixed and only heads are trained adversarially.
@@
where the instance weight $w_{\mu}^{(n)}$ is defined as 16. The nuisance $\mu$ is trained by maximizing the above risk, while the target parameter $\theta$ is trained by minimizing it. The other regularization terms in the overall loss function $\hat{J}(\theta,\mu)$ are the same as in 19. The pseudocode of NuSNet is presented in Appendix 0.C.

### v2

# A Distributionally Robust Framework for Nuisance in Causal Effect Estimation

@@ 1 Introduction
Causal inference enables us to assess the impact of a treatment action. Its application originated in the policy-making field [lalonde1986evaluating] including healthcare [sanchez2022causal]. Recently, the focus was expanded to individualized decision-making such as precision medicine [sanchez2022causal], recommendation [schnabel2016recommendations, bonner2018causal], and advertisement [sun2015causal, Wang2015RobustTC] with the help of advanced machine learning-based methods.
@@
We estimate the causal effect of a treatment action (e.g., prescription of a specific medicine). That is, we need accurate predictions of both potential outcomes with and without the treatment to take its difference since the supervision of the actual effect itself is never given. In observational data, actions are not assigned randomly but selected by past decision-makers. Therefore, models must generalize beyond the factual data distribution to handle the systematically missing counterfactual outcomes. This is called the fundamental problem of causal inference [shalit2017estimating].
@@
Conventional statistical methods for causal inference deal with this action selection bias by matching extraction or importance sampling [rosenbaum1983central]. A representative and versatile approach is inverse probability weighting using propensity scores (IPW) [austin2011introduction]. IPW estimates propensity scores—the probability of past treatment decisions—then weights instances by their inverse to account for selection bias. This two-step strategy has fundamental limitations: First, the final accuracy depends entirely on propensity score precision. Second, extreme weights can concentrate on few samples, reducing effective sample size and causing high estimation variance [kang2007demystifying]. The overall accuracy can only be guaranteed asymptotically, limiting its applicability to modern non-asymptotic scenarios such as high dimensional models as DNNs for capturing complex heterogeneity or complex action spaces.
@@
Various countermeasures have been tried to alleviate this problem, such as doubly robust (DR) methods [kang2007demystifying, kennedy2020towards, dudik2014doubly] and double machine learning [chernozhukov2018double, nie2021quasi], which are clever combinations of outcome prediction models and only weighting its residuals using estimated propensity. Nevertheless, the IPW-based approach’s limitation is the instability of the two-step procedure in non-asymptotic scenarios. Its large estimation variance is problematic in cases of high dimensional covariates or cases where the propensity score is close to 0 or $1$ and thus the effective sample size is limited [athey2018approximate].
@@
As in various other fields, advancement with deep neural networks (DNN) has gained substantial attention in causal inference literature [li2022survey]. One of the notable advancements made when [shalit2017estimating] and [johansson2016learning] introduced DNN to causal inference was representation-level balancing through distribution discrepancy measures. The representation extracted from the input covariates is encouraged to be balanced, i.e., independent of the action, by measuring and minimizing the discrepancy between the representation distributions conditioned on the action. Representation balancing provides non-asymptotic performance guarantees through generalization error bounds [shalit2017estimating]. This results in an end-to-end training procedure, free from the concerns in the intermediate estimation problem. However, it has been pointed out that the guarantee of generalization error of representation balancing requires the unrealistic assumption of invertible representation extraction [johansson2019support, zhao2019learning, wu2020representation]. In recent years, researchers have explored approaches that combine weighting with representation learning [Hassanpour2020Learning, wu2022learning, cheng2022learning]. However, to the best of our knowledge, this line of research has not yet yielded a theoretically grounded approach based on generalization error analysis.
@@
We analyze the generalization error of the weighted loss function using a propensity model, which is implemented as adversarial end-to-end learning of the target and weighting models. We define the worst-case loss with respect to the nuisance model ambiguity as a general framework. Our adversarial loss simultaneously accounts for the true propensity’s ambiguity and the statistical instability due to skewed weights. We demonstrate our framework’s effectiveness by applying it to two established methods: the doubly robust learner [kennedy2020towards] and representation-based architectures [curth2021nonparametric]. Experiments on both synthetic and real-world datasets show consistent improvements over baseline methods.
@@ 2 Problem setting
We consider a standard causal inference framework. We have observational data $D=\{(x^{(n)},a^{(n)},y^{(n)})\}_{1}^{N}$ with $N$ i.i.d. instances. Each instance contains a $d$-dimensional background feature $x^{(n)}\in\mathcal{X}\subset\mathbb{R}^{d}$, a treatment action $a^{(n)}\in\mathcal{A}=\{0,1\}$, and an outcome $y^{(n)}\in\mathcal{Y}$. In the Neyman-Rubin potential outcome framework [rubin2005causal], the potential outcomes of both factual and counterfactual actions are expressed as random variables $\{Y_{a}\}_{a\in\{0,1\}}$, of which only the factual outcome is observed $\left(y^{(n)}=y_{a^{(n)}}\right)$ and the counterfactual one $y_{1-a^{(n)}}$ is missing.
@@
As a sufficient condition for consistent learnability of the CATE, we follow the standard set of assumptions in the potential outcome framework [imbens2015causal].
@@ 3 Related work
Orthogonal statistical learning, such as the DR-Learner [kennedy2020towards] and the R-Learner [nie2021quasi], utilize two kinds of nuisance models of predicting outcomes and predicting treatments. These methods have been shown to be robust to estimation errors for the first-stage nuisance parameters of propensity and outcome models. Specifically, the errors do not affect the final estimation in the first-order sense of the Taylor expansion. However, their main limitation lies in high estimation variance when applied to non-asymptotic situations [athey2018approximate].
@@
Starting with [johansson2016learning, shalit2017estimating], a number of causal inference methods based on DNNs and representation balancing have been proposed [li2022survey]. The representation-based methods have been demonstrated to be superior in complex problems such as nonlinear responses [johansson2016learning], large actions spaces [tanimoto2021regret] including continuous [lopez2020cost] or structured spaces [harada2021graphite], and so forth. These are end-to-end methods based on adversarial formulations. They virtually evaluate the worst-case with respect to the uncertainty of the model by distribution discrepancy between the representations of covariates in treated and control groups. On the other hand, representation balancing has certain limitations in an unrealistic theoretical assumption that the representation extractor should be invertible. It is shown that unobservable error terms arise when the invertibility is violated [johansson2019support, zhao2019learning, wu2020representation].
A solution to this problem is the representation decomposition [Hassanpour2020Learning, wu2022learning, cheng2022learning]. They aim to identify confounding factors that affect both action selection and the outcomes and weights with only those factors. Joint optimization approaches have also been proposed for ATE estimation [shi2019adapting]. However, their formulations have no guarantee as a joint optimization, e.g., in the form of generalization error bound for all possible combinations of the target and nuisance models. Even though estimation of weights with other model parameters (i.e., outcome model and representations) fixed is guaranteed, and vice versa, though joint optimization may lead to unexpected results. Weights should be optimized to balance the distributions, but especially when the noise is heterogeneous, lowering the weights to the noisier regions may help the (wrongly) weighted loss of the outcome prediction model. Thus, we aim at a principled and versatile weighting method while incorporating the advantages of end-to-end modeling by adversarial formulation.
@@
Our proposed method can be interpreted in the context of distributionally robust optimization (DRO). DRO aims to achieve robust learning against discrepancies between the empirical distribution of training data and the actual test distribution [rahimian2019distributionally]. DRO is formalized as follows.
@@
Recent efforts in offline reinforcement learning revealed the benefit of pessimism on the candidate assessment [rashidinejad2021bridging, buckman2021importance]. In reinforcement learning, we consider modeling the cumulative expected reward in the long run as the Q function for assessing each action at each time step. The Q function is supposed to be maximized with respect to action $a$ during the inference phase. If the estimation error on an action is optimistic, i.e., if the Q value is overestimated, the action is likely to be selected over other better candidates. Therefore, conservative modeling of Q-function is preferable [kumar2020conservative], i.e., training a model to estimate below the true value when uncertain. The provable benefit of pessimism has been revealed in recent years [buckman2021importance]. We apply this pessimism principle to weighted estimation in causal inference; that is, our method pessimistically estimates the weighted loss with respect to the uncertainty of the weights. Our method applies the principle of pessimism in training that minimize the balanced loss with loss uncertainty due to propensity ambiguity and extreme weights, which can be evaluated by the generalization error analysis.
@@ 4.1 Adversarial Reformulation of Plug-in Nuisance
[…]
The assumed model classes are denoted by $M$ and $\Theta$. A typical form of ˜4 is instance weighted loss:
@@
Plug-in estimators typically analyze generalization error through stage-wise convergence rates [oprescu2019orthogonal]. For complex model classes, both convergence rates and their coefficients matter. We therefore adopt an end-to-end approach analyzing generalization error directly [shalit2017estimating], avoiding two-step estimation issues.
@@
The second term (b) in Section˜4.1 is upper-bounded by the excess risk of the weighted loss since $\mu^{*}\in M$.
$\text{(b)}\leq$ $L(\theta;\mu^{*})-\hat{L}(\theta;\mu^{*})$ (9)
@@
The proofs for these theorems can be found in Appendix˜0.A.
Consistent with findings and proposals in several prior studies [awasthi2024best, tanimoto2022improving, swaminathan2015counterfactual], our analysis also demonstrates that the mean squared weight is related to the stability of the loss (b) in Section˜4.1. In other words, assigning too large weights to a small fraction of the sample compromises the effective sample size. Therefore, to mitigate the impact of (b), we introduce a regularization on $\mu$ in the form of the squared weights, albeit at the potential cost of increasing the misspecification error (a) in Section˜4.1. More precisely, we define the class for restricted squared weights $\mathcal{R}$ as follows.
$\mathcal{R}=\left\{\mu\in M_{0}\middle|\frac{1}{N}\sum_{n}^{N}\left(w^{(n)}_{\mu}\right)^{2}\leq C\right\},$ (13)
@@ 4.2 Nuisance-robust Doubly-robust Network (NuDRNet)
Under the aforementioned approach, next, we will discuss the application to specific estimation methods. Among two-step methods with weighting, we take the doubly-robust learner or DRNet [kennedy2020towards] as a simple but clever baseline method. DRNet regresses a transformed target variable on $x$, which is calculated by the plug-in CATE estimate $\hat{f}_{1}(x^{(n)})-\hat{f}_{0}(x^{(n)})$ with residual adjusted with weights:
@@
We propose nuisance-robust DRNet (NuDRNet) as an extension of DRNet. Our approach starts with a pre-trained propensity $\hat{\mu}$ as the initial solution $\mu_{0}$ and a randomly initialized parameter $\theta$. We then perform adversarial optimization as described in ˜7, using the mean squared error: $\hat{L}(\theta;\mu)=\frac{1}{N}\sum_{n}\left(z^{(n)}_{\mu}-\tau_{\theta}(x^{(n)})\right)^{2}.$ Unlike DRNet, which uses a fixed pre-trained $\hat{\mu}$, NuDRNet perturbs $\hat{\mu}$ adversarially during training of $\hat{\tau}_{\theta}$.
Next, we discuss how to incorporate the constraint $\mu\in M$ into a gradient-based update. A typical implementation is as regularization term such as $\alpha\max\{0,\hat{E}(\mu)-c\},$ where $c$ is the tolerance parameter and is set to the evidence of the pre-trained solution with early stopping $c=\hat{E}(\mu_{0})$. However, since $\hat{L}(\theta;\mu)$ is not convex with respect to $\mu$, the regularization cannot reproduce the constrained optimization within $\mathcal{U}$. To address this issue, we implement the augmented Lagrangian method [bertsekas2014constrained] for handling the constraint $\mu\in\mathcal{U}$.
@@
In each epoch, we minimize $\hat{J}(\theta,\mu)$ with respect to $\theta$ and maximize with respect to $\mu$. Overall, our loss controls the error due to the uncertainty of $\mu$ (4.1-a) by maximizing the first term with respect to $\mu$ under the likelihood constraint in the third and the fourth terms, while simultaneously controlling the estimation variance of the weighted empirical loss (4.1-b) by flattening the weight with the second term. The parameters of evidence terms $\alpha_{k}$ and $\lambda_{k}$ are updated according to the augmented Lagrangian method. The whole algorithm is summarized in Algorithm˜1 and Figure˜1 illustrates the architecture.
@@
1: Train $f_{1},f_{0},\mu$ by an arbitrary supervised learning method, e.g.: $\hat{f}_{a}\leftarrow\operatorname*{arg\,min}_{f_{a}}\frac{1}{N}\sum_{n:a^{(n)}=a}(y^{(n)}-f_{a}(x^{(n)}))^{2}$ for each $a\in\{0,1\}$$\mu_{0}\leftarrow\operatorname*{arg\,min}_{\mu}-\frac{1}{N}\sum_{n}a^{(n)}\log\mu(x^{(n)})+(1-a^{(n)})\log(1-\mu(x^{(n)}))$
@@
6:   Update parameters with objective ˜19 and step sizes $\eta_{\theta},\eta_{\mu}$ from optimizes:
@@ 4.3 Nuisance-robust Shared Representation Network (NuSNet)
We further investigate exploiting representation-based architectures. Deep networks have also performed well in the context of causal inference owing to inductive bias of their representation layers. Among such methods, SNet [curth2021nonparametric] has shown outstanding performance thanks to its flexible architecture. SNet has three heads and five representation extractors. Specifically, there is a head for $y_{1}$, $y_{0}$, and $\mu$, respectively, and a confounder representation $\phi_{c}$ shared by the three heads, a representation shared by the two outcome heads $\phi_{o}$, and the representations unique to each hypothesis layer $\phi_{1},\phi_{0},\phi_{\mu}$. Denoting heads as $h_{\cdot}$, the SNet predictor is expressed as follows.
@@
The detailed architecture is illustrated in Figure˜1(a).
We apply adversarial joint learning to this architecture. The original SNet does not employ weighted loss. As noted before, simply applying non-adversarial joint optimization of weight and the outcome model leads to catastrophic results, as seen in our experiment in Appendix˜0.E. We therefore first pre-train as in the original SNet and then tune with an adversarially-weighted risk and other regularization terms. In the tuning phase, the shared representation layer is fixed and only heads are trained adversarially.
@@
where the instance weight $w_{\mu}^{(n)}$ is defined as ˜16. The nuisance $\mu$ is trained by maximizing the above risk, while the target parameter $\theta$ is trained by minimizing it. The other regularization terms in the overall loss function $\hat{J}(\theta,\mu)$ are the same as in ˜19. The pseudocode of NuSNet is presented in Appendix˜0.C.